find the domain of the function
g(x)= 9/9-4x
* note 9-4x is the bottom number of the fraction and 9 is the top number.

Answer by stanbon(75887)

find the domain of the function
g(x)= 9/(9-4x)
---
9-4x cannot be zero
---
9-4x = 0
4x = 9
x = 9/4
---------------
Domain: All Real Numbers except x = 9/4.
=============================================
Cheers,
Stan H.
============